OpenStudy (anonymous):

an inlet pipe can fill an empty pool in 5 hours, and another inlet pipe can fill the pool in 4 hours. how long will it take both pipes to fill the pool?

OpenStudy (anonymous):

First Inlet Pipe can fill the empty pool in ---------- 5 hours First Inlet Pipe's 1 hour work = 1/5 Similarly, Second Inlet Pipe's 1 hour work = 1/4 So, (First + Second) Inlet Pipe's work in 1 hour = (1/5 + 1/4) = 9/20 So, the time taken will be : 1/(9/20) = 20/9 Hours...

OpenStudy (anonymous):

x/5+x/4 = 1